OpenStudy (anonymous):

solve the equation by completing the square x^2-7x+1=0

OpenStudy (p0sitr0n):

\[x^2-7x+1=0\] \[(x^2-7x+12.25)-12.25+1=0\] \[(x-3.5)^2=11.25\] x-3.5 = \[\pm \sqrt{11.25}\] x =\[\pm \sqrt{11.25}\]+3.5